Solve for the variables usin algebra.
I have a parallelogram..the left hand angle is d degrees, the upper righht angle is c degrees, the bottom left angle is (b-10) degrees and the lower right angle is (b+10) degress, and I have to solve for variable b, c, and d.

Answer by scott8148(5883)

opposite angles are equal and consecutive angles are supplementary

(b+10)+(b-10)=180 __ 2b=180 __ b=90

c=b-10 __ c=80

d=b+10 __ d=100